# Unveiling Impact of Frequency Components on Membership Inference Attacks for Diffusion Models

### v1

@@ Abstract
Diffusion models have achieved tremendous success in image generation, but they also raise significant concerns regarding privacy and copyright issues. Membership Inference Attacks (MIAs) are designed to ascertain whether specific data were utilized during a model’s training phase. As current MIAs for diffusion models typically exploit the model’s image prediction ability, we formalize them into a unified general paradigm which computes the membership score for membership identification. Under this paradigm, we empirically find that existing attacks overlook the inherent deficiency in how diffusion models process high-frequency information. Consequently, this deficiency leads to member data with more high-frequency content being misclassified as hold-out data, and hold-out data with less high-frequency content tend to be misclassified as member data. Moreover, we theoretically demonstrate that this deficiency reduces the membership advantage of attacks, thereby interfering with the effective discrimination of member data and hold-out data. Based on this insight, we propose a plug-and-play high-frequency filter module to mitigate the adverse effects of the deficiency, which can be seamlessly integrated into any attacks within this general paradigm without additional time costs. Extensive experiments corroborate that this module significantly improves the performance of baseline attacks across different datasets and models.
@@ 1 Introduction
Diffusion models [35, 13] have achieved significant success in areas such as image generation [4, 33, 14] and video generation [40, 1], and have been widely applied in real life. However, this success has brought increasing attention to copyright and privacy issues from both academia and industry [38, 39, 45]. Recent research shows that diffusion models exhibit a strong memory effect regarding images in the training set, making the risk of privacy leakage a serious concern [39, 8].
@@
Membership Inference Attacks (MIAs) are crucial for assessing model privacy. Their core objective is to determine whether specific data were utilized during a model’s training phase [32]. Generally, MIAs exploit the overfitting characteristics of models. They achieve this by capturing the discrepancies in how the model fits the training data compared to other data, thereby enabling the execution of these attacks [42]. In the field of image generation, there has been extensive prior research on MIAs targeting Variational Autoencoders (VAEs) [12] and Generative Adversarial Networks (GANs) [3, 11]. However, due to the distinct training and generation mechanisms of diffusion models, these established attacks are mostly ineffective when applied to diffusion models [8].
@@
Recently, MIAs for diffusion models have gradually emerged as a research hotspot. Matsumoto et al. [24] proposed a query-based attack strategy that determines member data by analyzing the model’s loss function. However, this approach overlooks the uncertainty introduced by Gaussian noise. SecMI [8] is based on DDIM inversion [35, 15], obtaining intermediate outputs during generation through a deterministic inversion process. Nevertheless, this method demands a large number of queries, resulting in a significant increase in time costs. Kong et al. [16] introduced a proximal initialization method that utilizes model predictions to obtain initial noise. Although this approach somewhat reduces time overhead, the approximation of initial noise introduces errors that limit the attack performance. Zhai et al. [44] studied text-to-image diffusion models, exploring the associative memory between texts and images. In summary, the core mechanism of these methods is to quantify the model’s image recovery ability by calculating the pixel-wise error and use this as the basis for constructing the membership inference decision logic.
@@
From the perspective of frequency principles, diffusion models exhibit a distinctive generation process: they first denoise low-frequency signals representing overall structure and subsequently incorporate high-frequency details into the samples [41]. This fundamental processing asymmetry means diffusion models handle low-frequency components with greater fidelity and consistency, while high-frequency components show more variation in their reconstruction. While current membership inference attacks have significantly advanced the field, they have not explicitly considered how this frequency-dependent processing affects their effectiveness. This gap is important because a diffusion model’s varying behaviour across frequency bands directly influences its distinctive processing of training versus non-training images, the exact signal MIAs seek to detect. Our analysis within a general pixel-wise error-based MIA paradigm revealed that high-frequency components introduce substantial variance in membership scores, often leading to the misclassification of certain samples. We term this phenomenon "high-frequency deficiency".
@@
  • To the best of our knowledge, this study is the first to explore the impact of frequency domain information on MIAs targeting diffusion models. We formalize a general paradigm for existing pixel-wise error-based attacks and conduct an in-depth analysis of the impact of frequency domain information. The results reveal that existing attacks generally overlook the standard deviation induced by high-frequency deficiency, which restricts their attack performance.
  • To address this issue, we proposed a plug-and-play high-frequency filter module. This module effectively suppresses high-frequency deficiency, and we theoretically demonstrated its capacity to improve attack intensity. This module can be seamlessly integrated into all pixel-wise error-based attacks with negligible additional time overhead.
@@ 2 Related Works
Membership Inference Attacks. Shokri et al. [32] proposed the membership inference attacks, which primarily targeted classification models in machine learning. As the evolution of membership inference attacks continues, they can be classified into two categories: black-box attacks [31, 36, 5] and white-box attacks [25, 19], determined by the degree of access granted to the target model. In a white-box setting, the attacker can access the model parameters, whereas in a black-box setting, the attacker only receives the final output of the model. Moreover, noteworthy advancements have been achieved in membership inference attacks targeting generative models. Hayes et al. [11] demonstrated that membership can be effectively discerned through the logits of the discriminator in GANs. Hilprecht et al. [12] introduced a Monte Carlo scoring methodology incorporating the reconstruction loss term to facilitate attacks on VAEs.
Membership Inference Attacks on Diffusion Models. Recently, membership inference attacks on diffusion models have garnered increasing attention. In white-box settings, Pang et al. [27] proposed executing an attack through the utilization of gradient information extracted from loss. In grey-box settings, the attacker can only access the intermediate and final outputs [8]. Matsumoto et al. [24] pioneered the approach of employing diffusion loss to perform query-based membership inference. Duan et al. [8] introduced an attack leveraging DDIM inversion to retrieve intermediate outputs from the models. Meanwhile, Kong et al. [16] proposed a proximal initialization technique to acquire the deterministic initial noise. The attack is realized by the prediction of this noise from the model. In addition, attacks leveraging the correlation between texts and images have also achieved advancements [39, 44, 22]. Moreover, there has also been a growing interest in attacks for diffusion models in black-box settings [26, 20].
@@ 3 Preliminaries
Denoising Diffusion Implicit Model (DDIM). DDIM [35] upgrades the DDPM [13] framework by incorporating a non-Markovian process, which effectively decouples $x_{t-1}$ from $x_{t}$. This innovation allows for skipping timesteps, significantly accelerating the sampling process. DDIM redefines the denoising distribution as follows:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}\left(\sqrt{\overline{\alpha}_{t-1}}x_{0}+\sqrt{1-\overline{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\frac{x_{t}-\sqrt{\overline{\alpha}_{t}}x_{0}}{\sqrt{1-\overline{\alpha}_{t}}},\sigma_{t}^{2}I\right).$ (1)
where $\overline{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $(\alpha_{1},\ldots,\alpha_{T})$ are the predefined noise schedules, $\epsilon_{\theta}(x_{t},t)$ is predicted by the diffusion models, $\epsilon\sim\mathcal{N}(0,I)$. The denoising process defined by DDIM is outlined as follows:
@@
where $\sigma_{t}=\eta\sqrt{(1-\overline{\alpha}_{t-1})/(1-\overline{\alpha}_{t})}\sqrt{1-\overline{\alpha}_{t}/\overline{\alpha}_{t-1}}$ is the variance schedule, $\eta\in[0,1]$. The case $\eta=0$ corresponds to the DDIM, while $\eta=1$ corresponds to the deterministic DDPM. More details about the DDPM is provided in Appendix A.
@@ 4.1 Formalization of MIAs for Diffusion Models
Threat Model. Membership inference focuses on determining whether specific data were utilized in the training. Formally, consider a model $f_{\theta}$ parameterized by weights $\theta$ and a dataset $D=\{x_{1},\ldots,x_{n}\}$ sampled from data distribution $q_{data}$. Following established conventions [30, 2, 8], $D$ is split into two subsets, $D_{M}$ and $D_{H}$. $D_{M}$ is the member set of $f_{\theta}$ and $D_{H}$ is the hold-out set, such that $D=D_{M}\cup D_{H}$, $\varnothing=D_{M}\cap D_{H}$. So, $f_{\theta}$ is trained on $D_{M}$. Each sample $x_{i}$ is equipped with a membership identifier $m_{i}$, where $m_{i}=1$ if $x_{i}\sim D_{M}$; otherwise, $m_{i}=0$. The attackers have access to the $f_{\theta}$ and $D$ but lack knowledge of the specific partitioning between $D_{M}$ and $D_{H}$. The goal is to design an attack algorithm $\mathcal{A}$ that predicts the membership identifier $m_{i}$ for any given sample $x_{i}$:
@@
where $\mathcal{A}(x_{i},\theta)=1$ means $x_{i}$ comes from $D_{M}$, $\mathds{1}[A]=1$ if $A$ is true, and $\tau$ is the threshold. For generative models, we extend this framework by denoting the generator as $G_{\theta}$ with weights $\theta$ and the generative distribution as $p_{\theta}(x)$, where the generated samples $x\sim p_{\theta}(x)$.
General Paradigm. Attacks such as Naive [24], SecMI [8] and PIA [16], relying on pixel-wise errors, have demonstrated effectiveness in diffusion models. These attacks share common characteristics, utilizing the model’s image prediction capability to execute attacks. They can be unified as follows: given the image to be tested $x_{i}$, calculate the distance between the image $x_{i,t}$ predicted by the model at step $t$ and the target result $x_{i,t}^{target}$ of $x_{i}$ at step $t$, then using the distance as the membership score. $x_{i,t}$ and $x_{i,t}^{target}$ are obtained in different ways depending on the algorithm. Furthermore, setting a threshold $\tau$. If the score is less than $\tau$, the image is classified as member data; otherwise, it is classified as hold-out data. Formally, these attacks can be formulated as the general paradigm:
$\mathcal{A}(x_{i},\theta)=\mathds{1}\left[||x_{i,t}-x_{i,t}^{target}||_{q}\leq\tau\right],$ (6)
where $q$ represents the type of norm. We prove in Appendix B that the pixel-wise error-based attacks can be translated into the general paradigm expressed in Eq. 6.
@@ 4.2 Frequency Perspective of MIAs for Diffusion Models
Frequency Characteristics of Diffusion Models. Existing attack studies on diffusion models mainly concentrate on pixel-wise errors, yet they neglect an important aspect: analyzing models’ information processing from the frequency domain. Diffusion models’ operational mechanism features distinct frequency hierarchical properties. They first denoise low-frequency signals according to the learned distribution, then utilize specific low-frequency information as prior knowledge to process high-frequency details [28]. Recovering high-frequency information effectively relies not only on the model’s learned distribution but also closely ties to the image’s inherent structure and contours during denoising. Previous research [41] has shown that diffusion models exhibit more variation and uncertainty in handling high-frequency information.
Current pixel-wise error-based attacks mainly evaluate a model’s ability to process individual image. However, high- and low-frequency content varies greatly among images, and diffusion models have different mechanisms for handling such information. These two factors raise an interesting question: Does the high- and low-frequency content within a single image impact existing attack algorithms?
@@
Frequency Effects on MIAs for Diffusion. To study how frequency domain information affects MIAs, we visualized the relationship between high-frequency content and existing attacks’ membership scores. First, we transform images to the frequency domain using the Fourier transform and divide the high- and low-frequency regions by setting the frequency domain radius $=5$ as the high-low frequency boundary. Then, we calculated the percentage of high-frequency content by summing the squared frequency components in both regions. As shown in Fig. 1, visualizing the scores of attacks reveals a trend that as the high-frequency content of the images increased, so did the membership scores. This shows current attacks are biased, giving higher scores to images with more high-frequency content. Higher scores mean a lower degree of fitting, making it more likely to be classified as hold-out data. We term the phenomenon "high-frequency deficiency".
@@
Method STLU-10 Tiny-IN MS-COCO Flickr
@@ 4.3 High-Frequency Filter Design for Enhanced Diffusion MIAs
Based on the above observations, the following conclusion can be drawn: existing MIAs have not adequately accounted for the effect of high-frequency deficiency, which leads to their failure in certain specific scenarios. Intuitively, the inherent deficiency of diffusion models in handling high-frequency components leads to confusion in distinguishing between images with different high-frequency content. Further, we explore the effects of high-frequency deficiency from a theoretical perspective. Definition 1 gives an advantage measure that characterizes how well an algorithm can distinguish between member data and hold-out data. Theorem 1 characterizes the advantage of score-based MIAs for diffusion models.
@@
[42] The membership advantage of algorithm $\mathcal{A}$ is defined as:
$Adv^{M}(\mathcal{A})=\Pr[\mathcal{A}=1|m=1]-\Pr[\mathcal{A}=1|m=0].$ (7)
@@
Theorem 1.
Assume membership scores $s$ follow a normal distribution. Let membership scores of member data $s_{1}\sim N(0,\sigma_{M}^{2})$ and hold-out data $s_{2}\sim N(0,\sigma_{H}^{2})$, where $\sigma_{H}>\sigma_{M}$. The membership advantage of algorithm $\mathcal{A}$ is:
$Adv^{M}(\mathcal{A})=erf\left(\frac{\sigma_{H}}{\sigma_{M}}\sqrt{\frac{\ln(\sigma_{H}/\sigma_{M})}{(\sigma_{H}/\sigma_{M})^{2}-1}}\right)-erf\left(\sqrt{\frac{\ln(\sigma_{H}/\sigma_{M})}{(\sigma_{H}/\sigma_{M})^{2}-1}}\right),$ (8)
[…]
Proof. We provide the proof of Theorem 1 in Appendix C.
The advantage is 0 when $\sigma_{H}=\sigma_{M}$ and increases with the $\sigma_{H}/\sigma_{M}$. When $\sigma_{H}<\sigma_{M}$ means that there is no member advantage. Mathematically, the high-frequency deficiency contributes equally to a high standard deviation in the scores of both member and hold-out data, which leads to a decrease in $\sigma_{H}/\sigma_{M}$, weakening the advantages of identifying member data and consequently interfering with the effective differentiation between member and hold-out data.
@@
where $\odot$ denotes element-wise multiplication, and $\beta_{i,t}(r)$ is a mask designed as a filtering factor for frequency:
@@
$\mathcal{A}^{\prime}(x_{i},\theta)=\mathds{1}\left[||\mathcal{F}(x_{i,t})-\mathcal{F}(x_{i,t}^{target})||_{q}\leq\tau\right],$ (11)
@@
Denote the original standard deviations of membership scores in member and hold-out data as $\sigma_{M}$ and $\sigma_{H}$, and the standard deviations after removing the high-frequency components are $\sigma_{M}^{\prime}$ and $\sigma_{H}^{\prime}$. The standard deviation of membership scores in the high-frequency components is $h_{M}/h_{H}$, and the low-frequency components is $l_{M}/l_{H}$ in member and hold-out data. Let $l_{H}-l_{M}=\Delta$, $h_{M}=k\cdot h_{H}$ with $k>0$. If $k^{2}>1+\frac{2\Delta}{h_{H}^{2}}(l_{M}+2\Delta-\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}})$, we have:
@@
Proof Sketch.
Let $\sigma_{H}=\sigma_{H}^{\prime}+\Delta_{H}$ and $\sigma_{M}=\sigma_{M}^{\prime}+\Delta_{M}$, Eq. 12 is equivalent to:
$\Delta_{M}-\Delta_{H}>0.$ (13)
We can rewrite Eq. 13 in terms of $h_{H}$, $l_{M}$, $\Delta$ and $k$. We obtain:
$(k^{2}-1)h_{H}^{2}>2l_{M}\Delta+2\Delta^{2}-2\Delta\sqrt{(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}}.$ (14)
Some algebraic manipulation gets that:
$k^{2}>1+\frac{2\Delta}{h_{H}^{2}}(l_{M}+2\Delta-\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}}).$ (15)
We provide the detailed proof of Proposition 1 in Appendix D. ∎
Since $l_{M}+2\Delta-\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}}<0$, $\sigma_{H}^{\prime}/\sigma_{M}^{\prime}>\sigma_{H}/\sigma_{M}$ is hold constantly when $k\geq 1$. The key insight of Proposition 1 is that when the ratio of $h_{M}/h_{H}$ exceeds a certain threshold, filtering the membership scores derived from high-frequency components will amplify the membership advantage. It provides theoretical validation for our high-frequency filtering approach. To further validate the practical applicability of the theoretical framework, we systematically investigated the constraint conditions of $k$. As shown in Tab. 2, $k$ satisfies its constraint conditions under normal circumstances.
Method Tiny-IN STLU-10 MS-COCO Flickr
$k^{2}$ $f$ $k^{2}$ $f$ $k^{2}$ $f$ $k^{2}$ $f$
Naive 0.988 $>$ 0.921 0.976 $>$ 0.932 1.139 $>$ 0.924 0.910 $>$ 0.892
PIA 0.976 $>$ 0.927 1.290 $>$ 0.904 1.290 $>$ 0.914 1.317 $>$ 0.909
SecMI 1.102 $>$ 0.920 0.959 $>$ 0.897 1.778 $>$ 0.919 1.102 $>$ 0.911

### v2

@@ Abstract
Diffusion models have achieved tremendous success in image generation, but they also raise significant concerns regarding privacy and copyright issues. Membership Inference Attacks (MIAs) are designed to ascertain whether specific data were utilized during a model’s training phase. As current MIAs for diffusion models typically exploit the model’s image prediction ability, we formalize them into a unified general paradigm which computes the membership score for membership identification. Under this paradigm, we empirically find that existing attacks overlook the inherent deficiency in how diffusion models process high-frequency information. Consequently, this deficiency leads to member data with more high-frequency content being misclassified as hold-out data, and hold-out data with less high-frequency content tend to be misclassified as member data. Moreover, we theoretically demonstrate that this deficiency reduces the membership advantage of attacks, thereby interfering with the effective discrimination of member data and hold-out data. Based on this insight, we propose a plug-and-play high-frequency filter module to mitigate the adverse effects of the deficiency, which can be seamlessly integrated into any attacks within the general paradigm without additional time costs. Extensive experiments corroborate that this module significantly improves the performance of baseline attacks across different datasets and models.
@@ 1 Introduction
Diffusion models (Song et al., 2020a; Ho et al., 2020) have achieved significant success in areas such as image generation (Choi et al., 2021; Si et al., 2024; Huberman-Spiegelglas et al., 2024) and video generation (Wu et al., 2024; Blattmann et al., 2023), and have been widely applied in real life. However, this success has brought increasing attention to copyright and privacy issues from both academia and industry (Wang et al., 2023; Wen et al., 2024; Zhang et al., 2024). Recent research shows that diffusion models exhibit a strong memory effect regarding images in the training set, making the risk of privacy leakage a serious concern (Wen et al., 2024; Duan et al., 2023).
@@
Membership Inference Attacks (MIAs) are crucial for assessing model privacy. Their core objective is to determine whether specific data were utilized during a model’s training phase (Shokri et al., 2017). Generally, MIAs exploit the overfitting characteristics of models. They achieve this by capturing the discrepancies in how the model fits the training data compared to other data, thereby enabling the execution of these attacks (Yeom et al., 2018). In the field of image generation, there has been extensive prior research on MIAs targeting Variational Autoencoders (VAEs) (Hilprecht et al., 2019) and Generative Adversarial Networks (GANs) (Chen et al., 2020; Hayes et al., 2017). However, due to the distinct training and generation mechanisms of diffusion models, these established attacks are mostly ineffective when applied to diffusion models (Duan et al., 2023).
@@
Recently, MIAs for diffusion models have gradually emerged as a research hotspot. Matsumoto et al. (2023) proposed a query-based attack strategy that determines member data by analyzing the model’s loss function. SecMI (Duan et al., 2023) is based on DDIM inversion (Song et al., 2020a; Kim et al., 2022), obtaining intermediate outputs during generation through a deterministic inversion process. Kong et al. (2023) introduced a proximal initialization method that utilizes model predictions to obtain initial noise. Zhai et al. (2024) studied text-to-image diffusion models, exploring the associative memory between texts and images based on likelihood estimation. In summary, the core mechanism of these methods is to quantify the model’s image recovery ability and use this as the basis for constructing the membership inference decision logic.
@@
From the perspective of frequency principles, diffusion models exhibit a distinctive generation process: they first denoise low-frequency signals representing the overall structure and subsequently incorporate high-frequency details into the samples (Yang et al., 2023; Falck et al., 2025). This fundamental processing asymmetry means diffusion models handle low-frequency components with greater fidelity and consistency, while high-frequency components show more variation in their reconstruction. While current membership inference attacks have significantly advanced the field, they have not explicitly considered how this frequency-dependent processing affects their effectiveness. This gap is important because a diffusion model’s varying behaviour across frequency bands directly influences its distinctive processing of training versus non-training images, the exact signal MIAs seek to detect. Our analysis within a general error-based MIA paradigm revealed that high-frequency components introduce substantial standard deviation in membership scores, often leading to the misclassification of certain samples. We term this phenomenon "high-frequency deficiency".
@@
  • To the best of our knowledge, this study is the first to explore the impact of frequency domain information on MIAs targeting diffusion models. We formalize a general paradigm for existing error-based attacks and conduct an in-depth analysis of the impact of frequency domain information. The results reveal that existing attacks generally overlook the standard deviation induced by high-frequency deficiency, which restricts their attack performance.
  • To address this issue, we proposed a plug-and-play high-frequency filter module. This module effectively suppresses high-frequency deficiency, and we theoretically demonstrated its capacity to improve attack intensity. This module can be seamlessly integrated into all error-based attacks with negligible additional time overhead.
@@ 2 Related Works
Membership Inference Attacks. Shokri et al. (2017) proposed the membership inference attacks, which primarily targeted classification models in machine learning. As the evolution of membership inference attacks continues, they can be classified into two categories: black-box attacks (Salem et al., 2018; Song & Mittal, 2021; Choquette-Choo et al., 2021) and white-box attacks (Nasr et al., 2019; Leino & Fredrikson, 2020), determined by the degree of access granted to the target model. In a white-box setting, the attacker can access the model parameters, whereas in a black-box setting, the attacker only receives the final output of the model. Moreover, noteworthy advancements have been achieved in membership inference attacks targeting generative models. Hayes et al. (2017) demonstrated that membership can be effectively discerned through the logits of the discriminator in GANs. Hilprecht et al. (2019) introduced a Monte Carlo scoring methodology incorporating the reconstruction loss term to facilitate attacks on VAEs.
Membership Inference Attacks on Diffusion Models. Recently, membership inference attacks on diffusion models have garnered increasing attention. In white-box settings, Pang et al. (2023) proposed executing an attack through the utilization of gradient information extracted from loss. In grey-box settings, the attacker can only access the intermediate and final outputs (Duan et al., 2023). Matsumoto et al. (2023) pioneered the approach of employing diffusion loss to perform query-based membership inference. Duan et al. (2023) introduced an attack leveraging DDIM inversion to retrieve intermediate outputs from the models. Meanwhile, Kong et al. (2023) proposed a proximal initialization technique to acquire the deterministic initial noise. The attack is realized by the prediction of this noise from the model. In addition, attacks leveraging the correlation between texts and images have also achieved advancements (Wen et al., 2024; Zhai et al., 2024; Li et al., 2024). Moreover, there has also been a growing interest in attacks for diffusion models in black-box settings (Pang & Wang, 2023).
@@ 3 Preliminaries
Denoising Diffusion Implicit Model (DDIM). DDIM (Song et al., 2020a) upgrades the DDPM (Ho et al., 2020) framework by incorporating a non-Markovian process, which effectively decouples $x_{t-1}$ from $x_{t}$. This innovation allows for skipping timesteps, significantly accelerating the sampling process. DDIM redefines the denoising distribution as follows:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}\left(\sqrt{\overline{\alpha}_{t-1}}x_{0}+\sqrt{1-\overline{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\frac{x_{t}-\sqrt{\overline{\alpha}_{t}}x_{0}}{\sqrt{1-\overline{\alpha}_{t}}},\sigma_{t}^{2}I\right),$ (1)
where $\overline{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $(\alpha_{1},\ldots,\alpha_{T})$ are the predefined noise schedules, $\sigma_{t}$ is the variance schedule. Let $\epsilon_{\theta}(x_{t},t)$ denote the noise predicted by the denoising network. The denoising process is as follows:
@@
where $\epsilon\sim\mathcal{N}(0,I)$, $\sigma_{t}=\eta\sqrt{(1-\overline{\alpha}_{t-1})/(1-\overline{\alpha}_{t})}\sqrt{1-\overline{\alpha}_{t}/\overline{\alpha}_{t-1}}$, $\eta\in[0,1]$. The case $\eta=0$ corresponds to the deterministic DDIM, while $\eta=1$ corresponds to the DDPM. More details about the DDPM is provided in Appendix A.1.
@@ 4.1 Formalization of MIAs for Diffusion Models
Threat Model. Membership inference focuses on determining whether specific data were utilized in the training. Formally, consider a model $f_{\theta}$ parameterized by weights $\theta$ and a dataset $D=\{x_{1},\ldots,x_{n}\}$ sampled from data distribution $q_{data}$. Following established conventions (Sablayrolles et al., 2019; Carlini et al., 2022; Duan et al., 2023), $D$ is split into two subsets, $D_{M}$ and $D_{H}$. $D_{M}$ is the member set of $f_{\theta}$ and $D_{H}$ is the hold-out set, such that $D=D_{M}\cup D_{H}$, $\varnothing=D_{M}\cap D_{H}$. So, $f_{\theta}$ is trained on $D_{M}$. Each sample $x_{i}$ is equipped with a membership identifier $m_{i}$, where $m_{i}=1$ if $x_{i}\sim D_{M}$; otherwise, $m_{i}=0$. The attackers have access to the $f_{\theta}$ and $D$ but lack knowledge of the specific partitioning between $D_{M}$ and $D_{H}$. The goal is to design an attack algorithm $\mathcal{A}$ that predicts the membership identifier $m_{i}$ for any given sample $x_{i}$:
@@
where $\mathcal{A}(x_{i},\theta)=1$ means $x_{i}$ comes from $D_{M}$, $\mathds{1}[A]=1$ if $A$ is true, and $\tau$ is the threshold. For generative models, we extend this framework by denoting the generator as $G_{\theta}$ with weights $\theta$ and the generative distribution as $p_{\theta}(x)$, where the generated samples $x_{i}\sim p_{\theta}(x)$.
Adversary’s Capabilities. This work follows the widely used grey-box attack setting in prior studies (Duan et al., 2023; Matsumoto et al., 2023; Kong et al., 2023; Zhai et al., 2024). The grey-box setting is the mainstream attack setting for membership inference attacks against diffusion models. Given that diffusion models generate images through a progressive denoising process, the attacker is assumed to have access to the intermediate outputs produced during generation. This is achieved by specifying the timesteps at which these intermediate outputs are obtained, while access to the internal parameters of the model remains restricted.
General Paradigm. Attacks such as Naive (Matsumoto et al., 2023), SecMI (Duan et al., 2023) and PIA (Kong et al., 2023), relying on reconstruction errors, have demonstrated effectiveness in diffusion models. These attacks share common characteristics, utilizing the model’s image prediction capability to execute attacks. They can be unified as follows: given the image to be tested $x_{i}$, calculate the distance between the image $x_{i,t}$ predicted by the model at step $t$ and the target result $x_{i,t}^{target}$ of $x_{i}$ at step $t$, then using the distance as the membership score. $x_{i,t}$ and $x_{i,t}^{target}$ are obtained in different ways depending on the algorithm. Furthermore, setting a threshold $\tau$. If the score is less than $\tau$, the image is classified as member data; otherwise, it is classified as hold-out data. Formally, these attacks can be formulated as the general paradigm:
[…]
where $q$ represents the type of norm. We prove in Appendix B that the error-based attacks can be translated into the general paradigm expressed in Eq. 6.
@@ 4.2 Frequency Perspective of MIAs for Diffusion Models
Frequency Characteristics of Diffusion Models. Existing attack studies on diffusion models mainly concentrate on reconstruction errors, yet they neglect an important aspect: analyzing models’ information processing from the frequency domain. Diffusion models’ operational mechanism features distinct frequency hierarchical properties. They first denoise low-frequency signals according to the learned distribution, then utilize specific low-frequency information as prior knowledge to process high-frequency details (Qian et al., 2024). Recovering high-frequency information effectively relies not only on the model’s learned distribution but also closely ties to the image’s inherent structure and contours during denoising. Previous research (Yang et al., 2023) has shown that diffusion models exhibit more variation and uncertainty in handling high-frequency information.
Current error-based attacks mainly evaluate a model’s ability to process individual image. However, high- and low-frequency content varies greatly among images, and diffusion models have different mechanisms for handling such information. These two factors raise an interesting question: Does the high- and low-frequency content within a single image impact existing attack algorithms?
@@
Frequency Effects on MIAs for Diffusion. To study how frequency domain information affects MIAs, we visualized the relationship between high-frequency content and existing attacks’ membership scores. First, we transform images to the frequency domain using the Fourier transform and divide the high- and low-frequency regions by setting the frequency domain radius =5=5 as the high-low frequency boundary. Then, we calculated the percentage of high-frequency content by summing the squared frequency components in both regions. As shown in Fig. 1, visualizing the scores of attacks reveals a trend that as the high-frequency content of the images increased, so did the membership scores. This shows current attacks are biased, giving higher scores to images with more high-frequency content. Higher scores mean a lower degree of fitting, making it more likely to be classified as hold-out data. We term the phenomenon "high-frequency deficiency".
@@
Method
STLU-10 Tiny-IN MS-COCO Flickr
@@ 4.3 High-Frequency Filter Design for Enhanced Diffusion MIAs
Based on the above observations, the following conclusion can be drawn: existing MIAs have not adequately accounted for the effect of high-frequency deficiency, which leads to their failure in certain specific scenarios. Intuitively, the inherent deficiency of diffusion models in handling high-frequency components leads to confusion in distinguishing between images with different high-frequency content. Further, we explore the effects of high-frequency deficiency from a theoretical perspective. We rely on the previously established definition of membership inference attack capability (Yeom et al., 2018). Definition 1 gives an advantage measure that characterizes how well an algorithm can distinguish between member data and hold-out data.
@@
The membership advantage of algorithm $\mathcal{A}$ is defined as:
$Adv^{M}(\mathcal{A})=\Pr[\mathcal{A}=1|m=1]-\Pr[\mathcal{A}=1|m=0],$ (7)
@@
When membership scores follow a normal distribution, Yeom et al. (2018) reinterpret the membership advantage in terms of distributional differences. Specifically, they express the advantage using the ratio between the standard deviation of the hold-out scores $\sigma_{H}$ and that of the member scores $\sigma_{M}$. And they further demonstrate that this ratio is positively correlated with the membership advantage:
$\sigma_{H}/\sigma_{M}\propto Adv^{M}(\mathcal{A}).$ (8)
A larger ratio $\sigma_{H}/\sigma_{M}$ corresponds to a greater advantage for the attacker. When $\sigma_{H}<\sigma_{M}$, it means that there is no member advantage. Mathematically, the high-frequency deficiency contributes equally to a high standard deviation in the scores of both member and hold-out data, which may lead to a decrease in $\sigma_{H}/\sigma_{M}$, weakening the advantages of identifying member data and consequently interfering with the effective differentiation between member and hold-out data. This perspective is further validated in proposition 1.
@@
where $\odot$ denotes element-wise multiplication, and $\beta_{i,t}(r)$ is a mask designed as a filtering factor for frequency, it can be formalized as:
@@
[…]
@@
Denote the original standard deviations of membership scores in member and hold-out data as $\sigma_{M}$ and $\sigma_{H}$, and the standard deviations after removing the high-frequency components are $\sigma_{M}^{\prime}$ and $\sigma_{H}^{\prime}$. The standard deviation of membership scores in the high-frequency components is $h_{M}(h_{H})$, and the low-frequency components is $l_{M}(l_{H})$ in member(hold-out) data. Let $l_{H}-l_{M}=\Delta$, $h_{M}=k\cdot h_{H}$ with $k>0$. If $k^{2}>1+\frac{2\Delta}{h_{H}^{2}}(l_{M}+2\Delta-\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}})$, we have:
@@
Proof.
We provide the detailed proof of Proposition 1 in Appendix C. ∎
Since $l_{M}+2\Delta-\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}}<0$, $\sigma_{H}^{\prime}/\sigma_{M}^{\prime}>\sigma_{H}/\sigma_{M}$ is hold constantly when $k\geq 1$. The key insight of Proposition 1 is that when the ratio of $h_{M}/h_{H}$ exceeds a certain threshold, high-frequency contents will weaken the membership advantage, and filtering the membership scores derived from high-frequency components will amplify the membership advantage. From a theoretical perspective, it demonstrates that high-frequency components reduce the attack capability of the algorithm, while also validating the effectiveness of high-frequency filtering methods.
To further validate the practical applicability of the theoretical framework, we investigated the constraint conditions of $k$. The result shows that $k$ satisfies its constraint conditions under normal circumstances. We also provide a discussion on normality and verify the validity of the normality assumption in Proposition 1, while experimental results further demonstrate that our method does not rely on the assumption of normality. Detailed experiments and analysis are provided in Appendix D.2.